a) Define the surface tension and surface energy.

b) Explain the elastic limit with an example. Draw the stress-strain curve for elasticity.

c) A capillary tube with an inside radius of 638 π‘›π‘š can support a 100 π‘šπ‘š column of liquid that has a density of 820 π‘˜π‘”π‘šβˆ’3 . The observed contact angle is 12.5Β°. Find the surface tension of the liquid for β„Ž = 20 π‘π‘š.


1
Expert's answer
2021-12-07T09:54:03-0500

a) Surface tension: a property of liquid to take the least possible area. Surface energy: the physical quantity that defines how surface tension will work.

b) The elastic limit is the limit of mechanical extension of solids after which they do not experience plastic deformation.

c) Find the surface tension:


"\sigma=\frac{h\rho gr}{2\cos\theta}=5.25·10^{-4}\text{ N/m}."